# EXTENSIONS OF UNIRATIONAL GROUPS

### v2

@@ 3 Examples of extensions of unirational groups that are not unirational
We first note that $K/\mathbf{F}_{p}(\lambda,\mu)$ is separable, precisely because $\lambda,\mu$ are $p$-independent over $K$ [Mat, Th. 26.6]. Thus the non-unirationality of $U$ may be checked over $\mathbf{F}_{p}(\lambda,\mu)$ [Ros3, Th. 1.6], hence we may assume that $\lambda,\mu$ form a $p$-basis for $K$. We may also extend scalars to $K_{s}$ and thereby additionally assume that $K$ is separably closed. Assume for the sake of contradiction that $U$ is unirational. Then by Lemma 3.2, there exist $n>0$,
@@
and a homomorphism $g\colon V_{n,\alpha}\rightarrow U$ whose projection onto the $Y$ coordinate is nonzero. By [Ros1, Prop. 6.4], $g=(Y(S,(S_{j})_{j}),X_{0}(S,(S_{j})_{j}),\dots,X_{p-1}(S,(S_{j})_{j}))$, where
@@
Because the above expression has degree <p<p in $S$, it is an identity of $p$-polynomials in the variables $S,S_{j}$ [Ros1, Prop. 6.4]. We claim that
@@
We claim that $F$ is homogeneous of degree $p^{n-1}$. To prove this, we first extend scalars to $L_{1}:=K(\mu^{1/p})$. Taking $Z,Z_{j}$ (instead of $S,S_{j}$) to be the variables on $V_{1,\lambda}$, we have the map $h_{1}\colon U_{L_{1}}\rightarrow V_{1,\lambda}$ defined by the formula $(Y,(X_{i})_{i})\mapsto(Z,(Z_{j})_{j})$ where $Z:=X_{p-1}$, $Z_{j}:=X_{j}$ for 0<j<p10<j<p-1, and $Z_{0}:=X_{0}+\mu^{1/p}Y$. Because $\alpha\notin L_{1}^{p}=K^{p}(\mu)$, due to (3.9) and the equality $K^{p}(\lambda)\cap K^{p}(\mu)=K^{p}$, we may Lemma 3.3 to the map $h_{1}\circ g$, in conjunction with [Ros1, Prop. 6.4] to conclude that $(X_{0}+\mu^{1/p}Y)(S,(S_{j})_{j})$ is homogeneous of degree $p^{n-1}$ in the $S_{j}$ (by which we mean once one excludes the linear term involving $S$). Now we extend scalars to $L_{2}:=K((\mu/\lambda)^{1/p})$. Over $L_{2}$, we have the map $h_{2}\colon U_{L_{2}}\rightarrow V_{1,\lambda}$ defined by the formula $Z:=X_{p-1}$, $Z_{j}:=X_{j}$ for $j=0$ and 1<j<p11<j<p-1, and $Z_{1}:=X_{1}+(\mu/\lambda)^{1/p}Y$. (Here we again use $p>2$ to ensure that 1<p11<p-1.) Once more, we may apply Lemma 3.3 to $h_{2}\circ g$ to conclude that $Z_{0}=X_{0}$ is homogeneous of degree $p^{n-1}$ in the $S_{j}$. Since $X_{0}+\mu^{1/p}Y$ was as well, we deduce that $Y(S,(S_{j})_{j})$ is homogeneous of degree $p^{n-1}$ in those variables, hence $F$ is (see (3.5)). This proves the claim.
@@ 4 Ext-unirational groups
Ext-unirationality may be rephrased more canonically as follows. Given a smooth finite type $K$-group scheme $G$, let $G_{\mathrm{uni}}\subset G$ denote its maximal unirational $K$-subgroup scheme. Then $G_{\mathrm{uni}}$ is a normal subgroup of $G$ [Ros3, Cor. 7.11]. Now define a sequence of $K$-group quotients $G_{\mathrm{uni}}^{[n]}$ of $G$ recursively by the formulas
$G_{\operatorname{uni}}^{[0]}:=G,\hskip 21.68121ptG_{\operatorname{uni}}^{[n+1]}:=G_{\operatorname{uni}}^{[n]}/(G_{\operatorname{uni}}^{[n]})_{\operatorname{uni}},n\geq 0.$
Then $G$ is ext-unirational if and only if $G_{\operatorname{uni}}^{[n]}=1$ for some (equivalently, all) $n\geq 0$. This also shows that we may replace the condition in Definition 4.1 that each $G_{i}$ be normal in $G_{i+1}$ by the condition that it is normal in $G$ without altering the definition. Furthermore, the construction of the $G_{\operatorname{uni}}^{[n]}$ commutes with separable extension of the ground field [Ros3, Cor. 7.10], and as a consequence, ext-unirationality is insensitive to passage to separable extensions: If $L/K$ is a separable field extension, then $G$ is ext-unirational over $K$ if and only if it is so over $L$.
@@
We will discuss analogues of Corollary 4.4 in the next section.
@@
If $T\subset G$ is a torus in a connected linear algebraic $K$-group $G$, then $G$ is unirational if and only if $Z_{G}(T)$ is [Ros3, Prop. 7.12]. Applying this with $T$ a maximal torus, and using Lemma 2.3, one sees that unirationality of $G$ is equivalent to that of the unipotent group $Z_{G}(T)/T$, whence of its maximal wound quotient. The same statement holds true for ext-unirationality, thanks to the following analogous result.
@@ 5 Deducing unirationality from quotients
Corollary 4.4 says that, when $K$ has degree of imperfection $1$, unirationality of a group may be tested upon passage to a certain quotient of that group. In this section we will prove an analogue of this result for unipotent groups $U$ over fields of higher degree of imperfection. In particular, we will see that unirationality of $U^{\operatorname{\rm{ab}}}$ implies unirationality of $U$ when $K$ has degree of imperfection $\leq 2$. We will also show that this result is optimal by giving, over every field of degree of imperfection $\geq 3$, an example of a wound unipotent group $U$ such that $U^{\operatorname{\rm{ab}}}$ is unirational but $U$ is not (though $U$ must be ext-unirational by Proposition 4.3).
We first verify an analogous result for permawoundness in place of unirationality. The proof below makes reference to the notion of semiwoundness. Recall that a unipotent $K$-group scheme $U$ is called semiwound when it contains no copies of ${\mathbf{G}}_{\rm{a}}$ over $K$. This notion exhibits similar properties to the closely related notion of woundness, which is merely semiwoundness plus smoothness and connectedness; see [Ros2, Appendix A] for the basic properties of semiwound groups.
@@
All smooth unipotent groups are permawound over perfect fields [Ros2, Prop. 5.2(i)], so we may assume that $K$ is imperfect. The only if direction follows from the fact that permawoundness is inherited by quotients, so we now concentrate on the converse. Let us first assume that $U$ is commutative (which, though we do not know it yet, must be the case a posteriori when $U$ is wound, as all wound permawound unipotent groups are commutative). The group $U$ is killed by $p^{n}$ for some $n\geq 0$, and we proceed by induction on $n$, the $n=0$ case being trivial. So suppose that $n>0$. The map $p^{n-1}]\colon U\rightarrow U$ descends to a surjective map $U/[p]U\twoheadrightarrow[p^{n-1}]U$, hence $p^{n-1}]U$ is permawound. If we let $\overline{U}:=U/[p^{n-1}]U$, then $U/[p]U\twoheadrightarrow\overline{U}/[p]\overline{U}$, so the latter group is permawound, hence, by induction, so is $\overline{U}$, because it is killed by $p^{n-1}$. Thus $U$ an extension of the two permawound groups $\overline{U}$ and $p^{n-1}]U$, hence is itself permawound by Proposition 2.1.
Now let $U$ be a semiwound unipotent group such that $U^{\operatorname{\rm{ab}}}/[p]U^{\operatorname{\rm{ab}}}$ is permawound. We will show that $U$ is commutative, hence permawound by the already-treated commutative case. First we note that $U$ must be connected, as otherwise $U$ would admit a nontrivial étale unipotent quotient, hence a nontrivial commutative $p$-torsion étale quotient, which would imply that $U^{\operatorname{\rm{ab}}}/[p]U^{\operatorname{\rm{ab}}}$ is disconnected, in violation of [Ros2, Prop. 6.2]. If $U=1$ the assertion is immediate, so assume that $U\neq 1$. Then $U$ contains a nontrivial smooth connected central $K$-subgroup $U^{\prime}\subset U$. By dimension induction, the quotient $\overline{U}:=U/U^{\prime}$ is permawound, as $U^{\operatorname{\rm{ab}}}/[p]U^{\operatorname{\rm{ab}}}$ surjects onto $\overline{U}^{\operatorname{\rm{ab}}}/[p]\overline{U}^{\operatorname{\rm{ab}}}$. The commutator map $U^{2}\rightarrow U$ descends to a map $\overline{U}^{2}\rightarrow U$, which further descends to a map $\overline{U}_{w}^{2}\rightarrow U$, where $\overline{U}_{w}$ is the maximal wound quotient of $\overline{U}$. This map is constant by [Ros3, Cor. 10.3,Th. 1.9(ii)], hence $U$ is commutative, as claimed.
Finally, let $U$ be an arbitrary smooth unipotent $K$-group scheme such that $U^{\operatorname{\rm{ab}}}/[p]U^{\operatorname{\rm{ab}}}$ is permawound, and let $U_{s}\trianglelefteq U$ be the maximal split $K$-subgroup. Then $U/U_{s}$ is permawound by the semiwound case, hence so is $U$, by Proposition 2.1 and [Ros2, Prop. 5.3]. ∎
@@
Thanks to Proposition 5.1, we may assume that $U$ is commutative and $p$-torsion. When $K$ is perfect, the only wound group is the trivial group, and every smooth unipotent $K$-group is permawound [Ros2, Prop. 5.2(i)], so we may assume that $K$ is imperfect. If $U_{s}\trianglelefteq U$ denotes the maximal split $K$-subgroup, then $U$ is permawound if and only if $U/U_{s}$ is, by Proposition 2.1 and [Ros2, Prop. 5.3]. Since $U_{s}$ is killed under any $K_{s}$-homomorphism from $U$ into a wound group, we may assume $U$ is semiwound (that is, contains no copy of ${\mathbf{G}}_{\rm{a}}$). The only if direction now follows from [Ros3, Prop. 10.1], plus the connectedness of $U$ (to ensure woundness of the maximal wound quotient) [Ros2, Prop. 6.2].
Now suppose that $K$ has finite nonzero degree of imperfection, and assume that $U$ is not permawound. We will show that $\operatorname{Hom}_{K_{s}}(U,\mathscr{V})$ is infinite, where (by abuse of notation) $\mathscr{V}$ is any $K$-form of the $K_{s}$-group which is denoted by $\mathscr{V}$ in [Ros2, Def. 7.3]. By [Ros2, Th. 1.4], we have for some $n>0$ (>0>0 because $U$ is not permawound) an exact sequence
@@
with $W$ wound, permawound, commutative, and $p$-torsion. We will show that $\operatorname{Ext}^{1}_{K_{s}}(W,\mathscr{V})$ is finite but $\operatorname{Ext}^{1}_{K_{s}}({\mathbf{G}}_{\rm{a}},\mathscr{V})$ is infinite, which will imply that $\operatorname{Hom}_{K_{s}}(U,\mathscr{V})$ is infinite as well. We may assume that $K=K_{s}$ [Ros3, Cor. 10.5].
First we prove the finiteness of $\operatorname{Ext}^{1}(W,\mathscr{V})$. We have a homomorphism $\phi\colon\operatorname{Ext}^{1}(W,\mathscr{V})\rightarrow\operatorname{Hom}(W,\mathscr{V})$ defined as follows: Given an extension $E$ in the former group, the Verschiebung endomorphism of $E$ descends to a map $W\rightarrow\mathscr{V}$ from the quotient $W$ to the subgroup $\mathscr{V}$. Then $\ker(\phi)$ consists of the extensions with trivial Verschiebung, and this is a trivial group by [Ros2, Cor. 8.3]. Thus the desired finiteness follows from the finiteness of $\operatorname{Hom}(W,\mathscr{V})$.
@@
We may assume that $K$ is separably closed. If $U$ is not semiwound, then we may use the fact that $\mathrm{H}^{1}(X,{\mathbf{G}}_{\rm{a}})=0$ for every affine scheme $X$ and dimension induction to conclude. We therefore assume that $U$ is semiwound. If $K$ has infinite degree of imperfection, then the assertion is immediate, thanks to [Ros2, Prop. 6.3]. Thus we may assume that $K$ has finite degree of imperfection. Because permawound groups are unirational [Ros3, Th. 1.9], we may use the rigidity property of permawound groups [Ros2, Th. 1.5] and an easy induction to reduce to showing that any torsor for the $K$-group $V$ over a unirational $K$-scheme $X$ is also unirational, where $V$ is either $\operatorname{R}_{K^{1/p}/K}(\alpha_{p})$ or $\mathscr{V}$. But this assertion follows from [Ros3, Prop. 8.6]. ∎
@@
that sends a test scheme $T$ to the space of pointed $T$-morphisms $(X\times_{K}T,x_{T})\rightarrow(G\times_{K}T,1_{T})$. It turns out that it is represented by a smooth unipotent $K$-group scheme when $X$ is geometrically reduced of finite type and $G$ is wound unipotent, and in fact exhibits the somewhat stronger property that the corresponding scheme may be regarded as a subfunctor of the above morphism functor considered on the category of all $K$-schemes [Ros1, Th. 4.3].
@@
Assertion (ii) follows from (i), since $U$ admits a filtration by commutative $p$-torsion wound unipotent groups by [CGP, Prop. B.3.2], and every such group is of the form $\{F=0\}$ for some reduced $p$-polynomial $F$ [CGP, Prop. B.1.13]. (We may assume that $K$ is infinite, since otherwise $r=0$.) So we concentrate on (i). We may assume that $K$ is separably closed. By [Ros3, Lem. 5.4], one may write $D_{i}$ as a finite disjoint union of $K$-schemes of the form $\mathrm{Spec}(K(\alpha_{ij}^{1/p^{n_{ij}}}))$ with $\alpha_{ij}\in K$, hence $\operatorname{R}_{D_{i}/K}({\mathbf{G}}_{\rm{m}})\simeq\prod_{j}\operatorname{R}_{K(\alpha_{ij}^{1/p^{n_{ij}}})/K}({\mathbf{G}}_{\rm{m}})$, so we are reduced to the case in which $D_{i}=\mathrm{Spec}(K(\lambda_{i}^{1/p^{n_{i}}}))$ for all $i$.
@@
We proceed by induction on $r$, the case $r=0$ being trivial. Suppose that $r>0$. By the ubiquity property of permawound groups [Ros2, Th. 1.4], there is an exact sequence
@@
with $W$ permawound. We require a slightly more precise result, namely, that $m$ may be bounded depending only on $r$ and $d$. (Note that $p=\mathrm{char}(K)$ is determined by $d$.) This follows from the proof of ubiquity. In particular, $\dim(W)$ may be bounded depending only on $r$ and $d$. By the rigidity property of permawound groups [Ros2, Th. 1.5], together with the fact that the number of terms in a filtration as in that result is bounded in terms of $\dim(W)$ and $p$, we may in fact assume that $U$ is either $\operatorname{R}_{K^{1/p}/K}(\alpha_{p})$ or $\mathscr{V}$ (even though technically the former is not wound, being nonsmooth). In the former case, $U$ is totally nonsmooth, so the only map from a geometrically reduced $K$-scheme into $U$ is the 0 map. So assume that $U=\mathscr{V}$, and we must prove the existence of a suitable $C=C(p,r,n_{1},\dots,n_{r})$.
We may assume that all $n_{i}$ are as small as possible – that is, if $n_{i}>0$, then $\lambda_{i}\notin K^{p}$. If $n_{i}=0$, then $X_{1}\simeq\mathbf{A}^{1}$, so any map $\prod X_{i}\rightarrow U$ which vanishes on $\infty\times\prod_{i>1}X_{i}$ must be constant because $U$ is wound. Thus we may assume that $\lambda_{1}\notin K^{p}$ and $n_{1}>0$. By [Ros3, Lem. 5.3], there is a totally nonsmooth (over $K$) $K$-subgroup $N\subset M:=\mathscr{M}\mathit{or}((\mathbf{P}^{1}\backslash\lambda_{1}^{1/p^{n_{1}}},\infty),(\mathscr{V},0))^{+}$ such that $(M_{K(\lambda_{1}^{1/p^{\infty}})})_{s}\subset N_{K(\lambda_{1}^{1/p^{\infty}})}$, where, for a unipotent $L$-group $V$, $V_{s}$ denotes the maximal split $L$-subgroup of $V$.

### v3

@@ 3 Examples of extensions of unirational groups that are not unirational
We first note that $K/\mathbf{F}_{p}(\lambda,\mu)$ is separable, precisely because $\lambda,\mu$ are $p$-independent over $K$ [Mat, Th. 26.6]. Thus the non-unirationality of $U$ may be checked over $\mathbf{F}_{p}(\lambda,\mu)$ [Ros4, Th. 1.6], hence we may assume that $\lambda,\mu$ form a $p$-basis for $K$. We may also extend scalars to $K_{s}$ and thereby additionally assume that $K$ is separably closed. Assume for the sake of contradiction that $U$ is unirational. Then by Lemma 3.2, there exist $n>0$,
@@
and a homomorphism $g\colon V_{n,\alpha}\rightarrow U$ whose projection onto the $Y$ coordinate is nonzero. By [Ros2, Prop. 6.4], $g=(Y(S,(S_{j})_{j}),X_{0}(S,(S_{j})_{j}),\dots,X_{p-1}(S,(S_{j})_{j}))$, where
@@
Because the above expression has degree <p<p in $S$, it is an identity of $p$-polynomials in the variables $S,S_{j}$ [Ros2, Prop. 6.4]. We claim that
@@
We claim that $F$ is homogeneous of degree $p^{n-1}$. To prove this, we first extend scalars to $L_{1}:=K(\mu^{1/p})$. Taking $Z,Z_{j}$ (instead of $S,S_{j}$) to be the variables on $V_{1,\lambda}$, we have the map $h_{1}\colon U_{L_{1}}\rightarrow V_{1,\lambda}$ defined by the formula $(Y,(X_{i})_{i})\mapsto(Z,(Z_{j})_{j})$ where $Z:=X_{p-1}$, $Z_{j}:=X_{j}$ for 0<j<p10<j<p-1, and $Z_{0}:=X_{0}+\mu^{1/p}Y$. Because $\alpha\notin L_{1}^{p}=K^{p}(\mu)$, due to (3.9) and the equality $K^{p}(\lambda)\cap K^{p}(\mu)=K^{p}$, we may Lemma 3.3 to the map $h_{1}\circ g$, in conjunction with [Ros2, Prop. 6.4] to conclude that $(X_{0}+\mu^{1/p}Y)(S,(S_{j})_{j})$ is homogeneous of degree $p^{n-1}$ in the $S_{j}$ (by which we mean once one excludes the linear term involving $S$). Now we extend scalars to $L_{2}:=K((\mu/\lambda)^{1/p})$. Over $L_{2}$, we have the map $h_{2}\colon U_{L_{2}}\rightarrow V_{1,\lambda}$ defined by the formula $Z:=X_{p-1}$, $Z_{j}:=X_{j}$ for $j=0$ and 1<j<p11<j<p-1, and $Z_{1}:=X_{1}+(\mu/\lambda)^{1/p}Y$. (Here we again use $p>2$ to ensure that 1<p11<p-1.) Once more, we may apply Lemma 3.3 to $h_{2}\circ g$ to conclude that $Z_{0}=X_{0}$ is homogeneous of degree $p^{n-1}$ in the $S_{j}$. Since $X_{0}+\mu^{1/p}Y$ was as well, we deduce that $Y(S,(S_{j})_{j})$ is homogeneous of degree $p^{n-1}$ in those variables, hence $F$ is (see (3.5)). This proves the claim.
@@ 4 Ext-unirational groups
Ext-unirationality may be rephrased more canonically as follows. Given a smooth finite type $K$-group scheme $G$, let $G_{\mathrm{uni}}\subset G$ denote its maximal unirational $K$-subgroup scheme. Then $G_{\mathrm{uni}}$ is a normal subgroup of $G$ [Ros4, Cor. 7.11]. Now define a sequence of $K$-group quotients $G_{u}^{[n]}$ of $G$ recursively by the formulas
$G_{u}^{[0]}:=G,\hskip 21.68121ptG_{u}^{[n+1]}:=G_{u}^{[n]}/(G_{u}^{[n]})_{\operatorname{uni}},n\geq 0.$
Then $G$ is ext-unirational if and only if $G_{u}^{[n]}=1$ for some (equivalently, all sufficiently large) $n\geq 0$. This also shows that we may replace the condition in Definition 4.1 that each $G_{i}$ be normal in $G_{i+1}$ by the condition that it is normal in $G$ without altering the definition. Furthermore, the construction of the $G_{u}^{[n]}$ commutes with separable extension of the ground field [Ros4, Cor. 7.10], and as a consequence, ext-unirationality is insensitive to passage to separable extensions:
Proposition 4.2.
If $L/K$ is a separable field extension, then $G$ is ext-unirational over $K$ if and only if it is so over $L$.
The following lemma will be useful for proving results beyond the affine setting.
@@
We will discuss analogues of Corollary 4.5 over fields of higher degree of imperfection in the next section.
@@
If $T\subset G$ is a torus in a connected linear algebraic $K$-group $G$, then $G$ is unirational if and only if $Z_{G}(T)$ is [Ros4, Prop. 7.12]. Applying this with $T$ a maximal torus, and using Lemma 2.3, one sees that unirationality of $G$ is equivalent to that of the unipotent group $Z_{G}(T)/T$, whence of its maximal wound quotient. The same statement holds true for ext-unirationality, thanks to the following analogous result.
@@ 5 Deducing unirationality from quotients
Corollary 4.5 says that, when $K$ has degree of imperfection $1$, unirationality of a group may be tested upon passage to a certain quotient of that group. In this section we will prove an analogue of this result for unipotent groups $U$ over fields of higher degree of imperfection. In particular, we will see that unirationality of $U^{\operatorname{\rm{ab}}}$ implies unirationality of $U$ when $K$ has degree of imperfection $\leq 2$. We will also show that this result is optimal by giving, over every field of degree of imperfection $\geq 3$, an example of a wound unipotent group $U$ such that $U^{\operatorname{\rm{ab}}}$ is $p$-torsion and unirational even though $U$ fails to be unirational (though $U$ must be ext-unirational by Proposition 4.4).
We first verify an analogous result for permawoundness in place of unirationality. For the definition of permawoundness, see either [Ros3, Def. 1.2] or the beginning of §2. The proof below also makes reference to the notion of semiwoundness. Recall that a unipotent $K$-group scheme $U$ is called semiwound when it contains no copies of ${\mathbf{G}}_{\rm{a}}$ over $K$. This notion exhibits similar properties to the closely related notion of woundness, which is merely semiwoundness plus smoothness and connectedness; see [Ros3, Appendix A] for the basic properties of semiwound groups.
@@
All smooth unipotent groups are permawound over perfect fields [Ros3, Prop. 5.2(i)], so we may assume that $K$ is imperfect. The only if direction follows from the fact that permawoundness is inherited by quotients, so we now concentrate on the converse. Let us first assume that $U$ is commutative (which, though we do not know it yet, must be the case a posteriori when $U$ is wound, as all wound permawound unipotent groups are commutative). The group $U$ is killed by $p^{n}$ for some $n\geq 0$, and we proceed by induction on $n$, the $n=0$ case being trivial. So suppose that $n>0$. The map $p^{n-1}]\colon U\rightarrow U$ descends to a surjective map $U/[p]U\twoheadrightarrow[p^{n-1}]U$, hence $p^{n-1}]U$ is permawound. If we let $\overline{U}:=U/[p^{n-1}]U$, then $U/[p]U\twoheadrightarrow\overline{U}/[p]\overline{U}$, so the latter group is permawound, hence, by induction, so is $\overline{U}$, because it is killed by $p^{n-1}$. Thus $U$ an extension of the two permawound groups $\overline{U}$ and $p^{n-1}]U$, hence is itself permawound by Proposition 2.1.
Now let $U$ be a semiwound unipotent group such that $U^{\operatorname{\rm{ab}}}/[p]U^{\operatorname{\rm{ab}}}$ is permawound. We will show that $U$ is commutative, hence permawound by the already-treated commutative case. First we note that $U$ must be connected, as otherwise $U$ would admit a nontrivial étale unipotent quotient, hence a nontrivial commutative $p$-torsion étale quotient, which would imply that $U^{\operatorname{\rm{ab}}}/[p]U^{\operatorname{\rm{ab}}}$ is disconnected, in violation of [Ros3, Prop. 6.2]. If $U=1$ the assertion is immediate, so assume that $U\neq 1$. Then $U$ contains a nontrivial smooth connected central $K$-subgroup $U^{\prime}\subset U$. By dimension induction, the quotient $\overline{U}:=U/U^{\prime}$ is permawound, as $U^{\operatorname{\rm{ab}}}/[p]U^{\operatorname{\rm{ab}}}$ surjects onto $\overline{U}^{\operatorname{\rm{ab}}}/[p]\overline{U}^{\operatorname{\rm{ab}}}$. The commutator map $U^{2}\rightarrow U$ descends to a map $\overline{U}^{2}\rightarrow U$, which further descends to a map $\overline{U}_{w}^{2}\rightarrow U$, where $\overline{U}_{w}$ is the maximal wound quotient of $\overline{U}$. This map is constant by [Ros4, Cor. 10.3,Th. 1.9(ii)], hence $U$ is commutative, as claimed.
Finally, let $U$ be an arbitrary smooth unipotent $K$-group scheme such that $U^{\operatorname{\rm{ab}}}/[p]U^{\operatorname{\rm{ab}}}$ is permawound, and let $U_{s}\trianglelefteq U$ be the maximal split $K$-subgroup. Then $U/U_{s}$ is permawound by the semiwound case, hence so is $U$, by Proposition 2.1 and [Ros3, Prop. 5.3]. ∎
@@
Thanks to Proposition 5.1, we may assume that $U$ is commutative and $p$-torsion. When $K$ is perfect, the only wound group is the trivial group, and every smooth unipotent $K$-group is permawound [Ros3, Prop. 5.2(i)], so we may assume that $K$ is imperfect. If $U_{s}\trianglelefteq U$ denotes the maximal split $K$-subgroup, then $U$ is permawound if and only if $U/U_{s}$ is, by Proposition 2.1 and [Ros3, Prop. 5.3]. Since $U_{s}$ is killed under any $K_{s}$-homomorphism from $U$ into a wound group, we may assume $U$ is semiwound (that is, contains no copy of ${\mathbf{G}}_{\rm{a}}$). The only if direction now follows from [Ros4, Prop. 10.1], plus the connectedness of $U$ (to ensure woundness of the maximal wound quotient) [Ros3, Prop. 6.2].
Now suppose that $K$ has finite nonzero degree of imperfection, and assume that $U$ is not permawound. We will show that $\operatorname{Hom}_{K_{s}}(U,\mathscr{V})$ is infinite, where (by abuse of notation) $\mathscr{V}$ is any $K$-form of the $K_{s}$-group which is denoted by $\mathscr{V}$ in [Ros3, Def. 7.3]. By [Ros3, Th. 1.4], we have for some $n>0$ (>0>0 because $U$ is not permawound) an exact sequence
@@
with $W$ wound, permawound, commutative, and $p$-torsion. We will show that $\operatorname{Ext}^{1}_{K_{s}}(W,\mathscr{V})$ is finite but $\operatorname{Ext}^{1}_{K_{s}}({\mathbf{G}}_{\rm{a}},\mathscr{V})$ is infinite, which will imply that $\operatorname{Hom}_{K_{s}}(U,\mathscr{V})$ is infinite as well. We may assume that $K=K_{s}$ [Ros4, Cor. 10.5].
First we prove the finiteness of $\operatorname{Ext}^{1}(W,\mathscr{V})$. We have a homomorphism $\phi\colon\operatorname{Ext}^{1}(W,\mathscr{V})\rightarrow\operatorname{Hom}(W,\mathscr{V})$ defined as follows: Given an extension $E$ in the former group, the Verschiebung endomorphism of $E$ descends to a map $W\rightarrow\mathscr{V}$ from the quotient $W$ to the subgroup $\mathscr{V}$. Then $\ker(\phi)$ consists of the extensions with trivial Verschiebung, and this is a trivial group by [Ros3, Cor. 8.3]. Thus the desired finiteness follows from the finiteness of $\operatorname{Hom}(W,\mathscr{V})$.
@@
We may assume that $K$ is separably closed. If $U$ is not semiwound, then we may use the fact that $\mathrm{H}^{1}(X,{\mathbf{G}}_{\rm{a}})=0$ for every affine scheme $X$ and dimension induction to conclude. We therefore assume that $U$ is semiwound. If $K$ has infinite degree of imperfection, then the assertion is immediate, thanks to [Ros3, Prop. 6.3]. Thus we may assume that $K$ has finite degree of imperfection. Because permawound groups are unirational [Ros4, Th. 1.9], we may use the rigidity property of permawound groups [Ros3, Th. 1.5] and an easy induction to reduce to showing that any torsor for the $K$-group $V$ over a unirational $K$-scheme $X$ is also unirational, where $V$ is either $\operatorname{R}_{K^{1/p}/K}(\alpha_{p})$ or $\mathscr{V}$. But this assertion follows from [Ros4, Prop. 8.6]. ∎
@@
that sends a test scheme $T$ to the space of pointed $T$-morphisms $(X\times_{K}T,x_{T})\rightarrow(G\times_{K}T,1_{T})$. It turns out that it is represented by a smooth unipotent $K$-group scheme when $X$ is geometrically reduced of finite type and $G$ is wound unipotent, and in fact exhibits the somewhat stronger property that the corresponding scheme may be regarded as a subfunctor of the above morphism functor considered on the category of all $K$-schemes [Ros2, Th. 4.3].
@@
Assertion (ii) follows from (i), since $U$ admits a filtration by commutative $p$-torsion wound unipotent groups by [CGP, Prop. B.3.2], and every such group is of the form $\{F=0\}$ for some reduced $p$-polynomial $F$ [CGP, Prop. B.1.13]. (We may assume that $K$ is infinite, since otherwise $r=0$.) So we concentrate on (i). We may assume that $K$ is separably closed. By [Ros4, Lem. 5.4], one may write $D_{i}$ as a finite disjoint union of $K$-schemes of the form $\mathrm{Spec}(K(\alpha_{ij}^{1/p^{n_{ij}}}))$ with $\alpha_{ij}\in K$, hence $\operatorname{R}_{D_{i}/K}({\mathbf{G}}_{\rm{m}})\simeq\prod_{j}\operatorname{R}_{K(\alpha_{ij}^{1/p^{n_{ij}}})/K}({\mathbf{G}}_{\rm{m}})$, so we are reduced to the case in which $D_{i}=\mathrm{Spec}(K(\lambda_{i}^{1/p^{n_{i}}}))$ for all $i$.
@@
We proceed by induction on $r$, the case $r=0$ being trivial. Suppose that $r>0$. By the ubiquity property of permawound groups [Ros3, Th. 1.4], there is an exact sequence
@@
with $W$ permawound. We require a slightly more precise result, namely, that $m$ may be bounded depending only on $r$ and $d$. (Note that $p=\mathrm{char}(K)$ is determined by $d$.) This follows from the proof of ubiquity. In particular, $\dim(W)$ may be bounded depending only on $r$ and $d$. By the rigidity property of permawound groups [Ros3, Th. 1.5], together with the fact that the number of terms in a filtration as in that result is bounded in terms of $\dim(W)$ and $p$, we may in fact assume that $U$ is either $\operatorname{R}_{K^{1/p}/K}(\alpha_{p})$ or $\mathscr{V}$ (even though technically the former is not wound, being nonsmooth). In the former case, $U$ is totally nonsmooth, so the only map from a geometrically reduced $K$-scheme into $U$ is the 0 map. So assume that $U=\mathscr{V}$, and we must prove the existence of a suitable $C=C(p,r,n_{1},\dots,n_{r})$.
We may assume that all $n_{i}$ are as small as possible – that is, if $n_{i}>0$, then $\lambda_{i}\notin K^{p}$. If $n_{i}=0$, then $X_{1}\simeq\mathbf{A}^{1}$, so any map $\prod X_{i}\rightarrow U$ which vanishes on $\infty\times\prod_{i>1}X_{i}$ must be constant because $U$ is wound. Thus we may assume that $\lambda_{1}\notin K^{p}$ and $n_{1}>0$. By [Ros4, Lem. 5.3], there is a totally nonsmooth (over $K$) $K$-subgroup $N\subset M:=\mathscr{M}\mathit{or}((\mathbf{P}^{1}\backslash\lambda_{1}^{1/p^{n_{1}}},\infty),(\mathscr{V},0))^{+}$ such that $(M_{K(\lambda_{1}^{1/p^{\infty}})})_{s}\subset N_{K(\lambda_{1}^{1/p^{\infty}})}$, where, for a unipotent $L$-group $V$, $V_{s}$ denotes the maximal split $L$-subgroup of $V$.